# Reverse Diffusion Sequential Monte Carlo Samplers

### v1

@@ Appendix C Theoretical guarantees of RDSMC
We formally state the theorem providing sufficient conditions for asymptotic accuracy of RDSMC and unbiasedness of the estimate of the normalization constant. Our proof strategy for asymptotic accuracy is adapted from Wu et al. [30] and the proof for unbiasedness follows the techniques presented in Naesseth et al. [22].
@@
    the first marginal estimate $\hat{p}(x_{T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}),$ and the ratios of subsequent ratio $\hat{p}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})/\hat{p}(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})$ are positive and bounded,
@@
  • Assumption (a) can be satisfied by the clipping of the marginal estimates $\hat{p}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ within the range $[c,C]$ for some constants $c,C>0$.
@@
Theorem 3 (Chopin and Papaspiliopoulos [19] – Proposition 11.4).
@@
$w_{t}=\frac{\hat{p}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\,p(x_{t+1}\mid x_{t})}{\hat{p}(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})\,q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})}.$
@@
$\log w_{t}=\log\frac{\hat{p}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})}{\hat{p}(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})}+\log\frac{p(x_{t+1}\mid x_{t})}{q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})}.$
The boundedness of $\log\frac{\hat{p}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})}{\hat{p}(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})}$ is guaranteed by Assumption (a). We now show the boundedness of $\log\frac{p(x_{t+1}\mid x_{t})}{q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})}$. First, note that
$p(x_{t+1}\mid x_{t})=\mathcal{N}(x_{t+1}\mid x_{t}+f_{t}x_{t}\delta,g_{t}^{2}\delta),$
@@
$\hat{Z}=\prod_{t=0}^{T}\frac{1}{N}\sum_{i=1}^{N}w_{t}^{(i)},\quad\text{where }w_{t}^{(i)}=\frac{\hat{p}_{t}^{(i)}p(x_{t+1}^{(i)}\mid x_{t}^{a^{i}_{t}})}{\hat{p}_{t+1}^{(i)}q(x_{t}^{a^{i}_{t}}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}^{(i)}})},\ w_{T}^{(i)}=\frac{\hat{p}_{T}^{(i)}}{\hat{p}(x_{T}^{a^{i}_{t}})}.$
@@
$=p(x_{T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}})p(x_{0:{T-1}},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0:{T-1}}},a_{0:T})$
$=p(x_{T}^{a_{T}},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}^{a_{T}})p(x_{T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}\mid x_{T}^{a_{T}},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}^{a_{T}})\prod_{t=0}^{T-1}p(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t}^{a_{t}},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}^{a_{t}})q(x_{t}^{a_{t}}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}^{a_{t}}\mid x_{t}^{a_{t}}),$
$=p(x_{T}^{a_{T}})\Big{(}\prod_{i=1}^{N}\frac{w_{T}^{(i)}}{\sum_{i=1}^{N}w_{T}^{(i)}}\Big{)}\prod_{i=1}^{N}\Big{[}\Big{(}\prod_{t=0}^{T-1}\frac{w_{t}^{(i)}}{\sum_{i=1}^{N}w_{t}^{(i)}}q(x_{t}^{a_{t}^{i}}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}^{(i)}})\Big{)}\Big{]}$
@@
By $w_{t}^{(i)}=\frac{\hat{p}_{t}^{(i)}p(x_{t+1}^{(i)}|x_{t}^{a_{t}^{i}})}{\hat{p}_{t+1}^{(i)}q(x_{t}^{a_{t}^{i}}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}^{(i)}})}$ and $w_{T}^{(i)}=\frac{\hat{p}_{T}^{(i)}}{p(x_{T}^{a_{T}^{i}})}$, we have:
@@
$=\Big{(}\prod_{t=0}^{T}\sum_{i=1}^{N}w_{t}^{(i)}\Big{)}^{-1}\hat{p}_{0}^{(1)}\Big{(}\prod_{t=0}^{T-1}p(x_{t+1}^{(1)}|x_{t}^{a_{t}^{1}})\Big{)}$
$\prod_{i=2}^{N}\hat{p}_{T}^{(i)}\prod_{i=2}^{N}\Big{[}\Big{(}\prod_{t=0}^{T-1}\widetilde{w}_{t}^{(i)}q(x_{t}^{a_{t}^{i}}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}^{(i)}})\Big{)}\Big{]}\Big{[}q({\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}^{a_{T}}\mid x_{T}^{a_{T}})\prod_{t=0}^{T-1}q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}^{a_{t}}\mid x_{t}^{a_{t}})\Big{]},$
@@
$\hat{Z}Q(x_{0:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0:T}},a_{0:T})$ $=\hat{Z}\Big{(}\prod_{t=0}^{T}\sum_{i=1}^{N}w_{t}^{(i)}\Big{)}^{-1}\hat{p}_{0}^{(1)}\Big{(}\prod_{t=0}^{T-1}p(x_{t+1}^{(1)}|x_{t}^{a_{t}^{1}})\Big{)}$
$\prod_{i=2}^{N}\Big{[}\hat{p}_{T}^{(i)}\Big{(}\prod_{t=0}^{T-1}\widetilde{w}_{t}^{(i)}q(x_{t}^{a_{t}^{i}}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}^{(i)}})\Big{)}\Big{]}\Big{[}q({\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}^{a_{T}}\mid x_{T}^{a_{T}})\prod_{t=0}^{T-1}q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}^{a_{t}}\mid x_{t}^{a_{t}})\Big{]}$
$=N^{-T}\hat{p}_{0}^{(1)}q(u_{0}^{a_{0}^{1}}\mid x_{0}^{a_{0}^{1}})\Big{(}\prod_{t=0}^{T-1}p(x_{t+1}^{(1)}|x_{t}^{a_{t}^{1}})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}^{a_{t+1}^{1}}\mid x_{t+1}^{a_{t+1}^{1}})\Big{)}$
$\prod_{i=2}^{N}\Big{[}\hat{p}_{T}^{(i)}q(u_{t}^{a^{i}_{T}}\mid x_{t}^{a^{i}_{T}})\Big{(}\prod_{t=0}^{T-1}\widetilde{w}_{t}^{(i)}q(x_{t}^{a_{t}^{i}}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}^{(i)}})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}^{a_{t+1}^{(i)}}\mid x_{t+1}^{a_{t+1}^{(i)}})\Big{)}\Big{]},$
@@
$\prod_{i=2}^{N}\Big{[}\hat{p}_{T}^{(i)}q(u_{t}^{a^{i}_{T}}\mid x_{t}^{a^{i}_{T}})\Big{(}\prod_{t=0}^{T-1}\widetilde{w}_{t}^{(i)}q(x_{t}^{a_{t}^{i}}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}^{(i)}})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}^{a_{t+1}^{(i)}}\mid x_{t+1}^{a_{t+1}^{(i)}})\Big{)}\Big{]}dx_{0:T}d{\color[rgb]{0.3125,0.46875,0.62890625}u_{0:T}}$
$=$ $N^{-T}\sum_{{a_{0:T}^{1}}}\int\int\hat{p}_{0}^{(1)}q(u_{0}^{a_{0}^{1}}\mid x_{0}^{a_{0}^{1}})\Big{(}\prod_{t=0}^{T-1}p(x_{t+1}^{(1)}|x_{t}^{a_{t}^{1}})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}^{a_{t+1}^{1}}\mid x_{t+1}^{a_{t+1}^{1}})\Big{)}dx_{0:T}^{1}du_{0:T}^{1}$
@@
Appendix D Experiment details
This section provides additional experiment details. §˜D.1 describes the implementation and hyperparameter details of RDSMC and all competing baselines. §˜D.2, D.3 and D.4 include supplementary information and additional results for different targets in the main experiments.

### v2

@@ Appendix C Theoretical guarantees of RDSMC
We formally state the theorem providing sufficient conditions for asymptotic accuracy of RDSMC and unbiasedness of the estimate of the normalization constant. Our proof strategy for asymptotic accuracy is adapted from wu2023practical and the proof for unbiasedness follows the techniques presented in naesseth2019elements.
@@
    the first marginal estimate $\hat{\pi}(x_{T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}),$ and the ratios of subsequent ratio $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})/\hat{\pi}(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})$ are positive and bounded,
@@
  • Assumption (a) can be satisfied by the clipping of the marginal estimates $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ within the range $[c,C]$ for some constants $c,C>0$.
@@
Theorem 3 (chopin2020introduction – Proposition 11.4).
@@
$w_{t}=\frac{\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\,\pi(x_{t+1}\mid x_{t})}{\hat{\pi}(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})\,q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})}.$
@@
$\log w_{t}=\log\frac{\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})}{\hat{\pi}(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})}+\log\frac{\pi(x_{t+1}\mid x_{t})}{q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})}.$
The boundedness of $\log\frac{\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})}{\hat{\pi}(x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})}$ is guaranteed by Assumption (a). We now show the boundedness of $\log\frac{\pi(x_{t+1}\mid x_{t})}{q(x_{t}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})}$. First, note that
$\pi(x_{t+1}\mid x_{t})=\mathcal{N}(x_{t+1}\mid x_{t}+f_{t}x_{t}\delta,g_{t}^{2}\delta),$
@@
$\hat{Z}=\prod_{t=0}^{T}\frac{1}{N}\sum_{i=1}^{N}w_{t}^{(i)},\quad\text{where }w_{t}^{(i)}=\frac{\hat{\pi}_{t}^{(i)}\pi(x_{t+1}^{(i)}\mid x_{t}^{a^{i}_{t}})}{\hat{\pi}_{t+1}^{(i)}q(x_{t}^{a^{i}_{t}}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}^{(i)}})},\ w_{T}^{(i)}=\frac{\hat{\pi}_{T}^{(i)}}{\hat{\pi}(x_{T}^{a^{i}_{t}})}.$
@@
$=\pi(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}})p(x_{0:{T-1}},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0:{T-1}}},a_{0:T})$
$=\pi(x_{t}^{a_{T}},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}^{a_{T}})\pi(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}\mid x_{T}^{a_{T}},{\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}^{a_{T}})\prod_{t=0}^{T-1}p(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}\mid x_{t}^{a_{t}},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}^{a_{t}})q(x_{t}^{a_{t}}\mid x_{t+1},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}^{a_{t}}\mid x_{t}^{a_{t}}),$
$=\pi(x_{t}^{a_{T}})\Big(\prod_{i=1}^{N}\frac{w_{T}^{(i)}}{\sum_{i=1}^{N}w_{T}^{(i)}}\Big)\prod_{i=1}^{N}\Big[\Big(\prod_{t=0}^{T-1}\frac{w_{t}^{(i)}}{\sum_{i=1}^{N}w_{t}^{(i)}}q(x_{t}^{a_{t}^{i}}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}^{(i)}})\Big)\Big]$
@@
By $w_{t}^{(i)}=\frac{\hat{\pi}_{t}^{(i)}\pi(x_{t+1}^{(i)}|x_{t}^{a_{t}^{i}})}{\hat{\pi}_{t+1}^{(i)}q(x_{t}^{a_{t}^{i}}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}^{(i)}})}$ and $w_{T}^{(i)}=\frac{\hat{\pi}_{T}^{(i)}}{q(x_{T}^{a_{T}^{i}})}$, we have:
@@
$=\Big(\prod_{t=0}^{T}\sum_{i=1}^{N}w_{t}^{(i)}\Big)^{-1}\hat{\pi}_{0}^{(1)}\Big(\prod_{t=0}^{T-1}\pi(x_{t+1}^{(1)}|x_{t}^{a_{t}^{1}})\Big)$
$\prod_{i=2}^{N}\hat{\pi}_{T}^{(i)}\prod_{i=2}^{N}\Big[\Big(\prod_{t=0}^{T-1}\widetilde{w}_{t}^{(i)}q(x_{t}^{a_{t}^{i}}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}^{(i)}})\Big)\Big]\Big[q({\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}^{a_{T}}\mid x_{T}^{a_{T}})\prod_{t=0}^{T-1}q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}^{a_{t}}\mid x_{t}^{a_{t}})\Big],$
@@
$\hat{Z}Q(x_{0:T},{\color[rgb]{0.3125,0.46875,0.62890625}u_{0:T}},a_{0:T})$ $=\hat{Z}\Big(\prod_{t=0}^{T}\sum_{i=1}^{N}w_{t}^{(i)}\Big)^{-1}\hat{\pi}_{0}^{(1)}\Big(\prod_{t=0}^{T-1}\pi(x_{t+1}^{(1)}|x_{t}^{a_{t}^{1}})\Big)$
$\prod_{i=2}^{N}\Big[\hat{\pi}_{T}^{(i)}\Big(\prod_{t=0}^{T-1}\widetilde{w}_{t}^{(i)}q(x_{t}^{a_{t}^{i}}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}^{(i)}})\Big)\Big]\Big[q({\color[rgb]{0.3125,0.46875,0.62890625}u_{T}}^{a_{T}}\mid x_{T}^{a_{T}})\prod_{t=0}^{T-1}q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t}}^{a_{t}}\mid x_{t}^{a_{t}})\Big]$
$=N^{-T}\hat{\pi}_{0}^{(1)}q(u_{0}^{a_{0}^{1}}\mid x_{0}^{a_{0}^{1}})\Big(\prod_{t=0}^{T-1}\pi(x_{t+1}^{(1)}|x_{t}^{a_{t}^{1}})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}^{a_{t+1}^{1}}\mid x_{t+1}^{a_{t+1}^{1}})\Big)$
$\prod_{i=2}^{N}\Big[\hat{\pi}_{T}^{(i)}q(u_{t}^{a^{i}_{T}}\mid x_{t}^{a^{i}_{T}})\Big(\prod_{t=0}^{T-1}\widetilde{w}_{t}^{(i)}q(x_{t}^{a_{t}^{i}}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}^{(i)}})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}^{a_{t+1}^{(i)}}\mid x_{t+1}^{a_{t+1}^{(i)}})\Big)\Big],$
@@
$\prod_{i=2}^{N}\Big[\hat{\pi}_{T}^{(i)}q(u_{t}^{a^{i}_{T}}\mid x_{t}^{a^{i}_{T}})\Big(\prod_{t=0}^{T-1}\widetilde{w}_{t}^{(i)}q(x_{t}^{a_{t}^{i}}\mid x_{t+1}^{(i)},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}^{(i)}})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}^{a_{t+1}^{(i)}}\mid x_{t+1}^{a_{t+1}^{(i)}})\Big)\Big]dx_{0:T}d{\color[rgb]{0.3125,0.46875,0.62890625}u_{0:T}}$
$=$ $N^{-T}\sum_{{a_{0:T}^{1}}}\int\int\hat{\pi}_{0}^{(1)}q(u_{0}^{a_{0}^{1}}\mid x_{0}^{a_{0}^{1}})\Big(\prod_{t=0}^{T-1}\pi(x_{t+1}^{(1)}|x_{t}^{a_{t}^{1}})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t+1}}^{a_{t+1}^{1}}\mid x_{t+1}^{a_{t+1}^{1}})\Big)dx_{0:T}^{1}du_{0:T}^{1}$
@@
Appendix D Comparison to Twisted Diffusion Sampler
In this section, we discuss similarities and differences between RDSMC and Twisted Diffusion Sampler [TDS, wu2023practical], which is an SMC method for conditional generation from diffusion models.
Structurally, our method shares similarities with TDS, as both methods use some form of "look-ahead" or "twisting" functions [naesseth2019elements, Chapter 3] to improve the efficiency of SMC. These twisting functions introduce future information to intermediate states, promote promising particles through the weighting procedure, and improve the SMC sampling efficiency.
Moreover, the incremental weight expression of both TDS and RDSMC takes the same form (their Eq. (11) v.s. our Eq.˜15):
$w_{t}$ $:=\frac{\psi(x_{t})\gamma^{0}_{t}(x_{t:T})}{\psi(x_{t+1})\gamma^{0}_{t}(x_{t+1:T})q(x_{t}\mid x_{t+1})},$
where $\psi$ is the twisting function, $\gamma^{0}_{t}$ is the untwisted intermediate target and $q$ is the proposal. The product $\gamma_{t}(x_{t:T})\coloneqq\psi(x_{t})\gamma_{t}^{0}(x_{t:T})$ defines a series of twisted intermediate targets.
While the expressions of these components are method-specific, the structure of the weighting functions is shared between TDS our RDSMC and follows from the generic twisted SMC framework [naesseth2019elements, Chapter 3].
Importantly, this weight construction introduces twisting functions in a telescoping manner such that their approximation errors cancel out over the full trajectory. This property ensures that the final target is correct, and the resulting SMC procedure is asymptotically exact for both TDS and RDSMC.
However, we want to highlight several important differences between TDS and RDSMC:
  1. 1.
    Problem setting: The problem we address is fundamentally different from that of TDS. TDS is designed to sample from the conditional distribution $p_{\theta}(x_{0}\mid y)\propto p_{\theta}(x_{0})p(y\mid x_{0})$ given a pretrained diffusion model $p_{\theta}(x_{0})$ and a likelihood model $p(y\mid x_{0})$ with the observation $y$. As a comparison, our goal is to sample from an arbitrary distribution $\pi(x_{0})$, assuming access to its unnormalized density $\tilde{\pi}(x_{0})$.
  2. 2.
    Design of key components: Due to the difference in settings, our intermediate target distributions, proposal distributions, and twisting functions require different designs. In particular,
    • TDS’s twisting function is $p(y\mid\hat{x}_{\theta}(x_{t})$ (their Eq. (8)), the likelihood evaluated at the denoising prediction $\hat{x}_{\theta}(x_{t})$ by the pretrained model, which approximates the intractable likelihood $p_{\theta}(y\mid x_{t})$. This twisting function incorporates the observation $y$ into the weighting function (their Eq. (11)). Notably, when $t=0$, TDS’s twisting function recovers the exact likelihood $p(y\mid x_{0})$ and no approximation is involved.
      Given the twisting function, TDS’s intermediate target is $p\left(y\mid\hat{x}_{\theta}(x_{t})\right)p_{\theta}(x_{t:T})$ where $p_{\theta}(x_{t:T})$ is the distribution under the pretrained diffusion model. The proposal is obtained by adding a term that is the gradient of the twisting function, $\nabla_{x_{t}}\log p(y\mid\hat{x}_{\theta}(x_{t}))$, to the pretrained diffusion model’s transition kernel (see their Eq (9) and (11)).
    • In our case, RDSMC uses the marginal estimate $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$ (Eq.˜10) as the twisting function, which approximates the exact marginal $\pi(x_{t})$. The marginal integrates out the future information in $\pi(x_{0})$, thereby encoding the target information into the intermediate weight.
      We define intermediate targets as $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\pi(x_{t:T})q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}}\mid x_{t:T})$ (a restatement of Eq.˜11); the product term $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})\pi(x_{t:T})$ is structurally similar to TDS’s intermediate target, while we additionally account for MC estimation randomness $q({\color[rgb]{0.3125,0.46875,0.62890625}u_{t:T}}\mid x_{t:T})$. We also utilize twisting function to inform the proposal, that is, we use the score estimate $s(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$, which is connected to the twisting function / marginal estimate $\hat{\pi}(x_{t},{\color[rgb]{0.3125,0.46875,0.62890625}u_{t}})$, to design the proposal in Eq.˜8.
  3. 3.
    Use of auxiliary variables: Our method introduces auxiliary randomness through the Monte Carlo estimation of the twisting function. Incorporating this randomness places RDSMC in the category of nested SMC methods [naesseth2015nested]. In contrast, TDS does not use auxiliary variables and operates more like a standard SMC algorithm, without requiring nested sampling.
Appendix E Experiment details
This section provides experimental details and additional results.
§˜E.1 describes the implementation and hyperparameter settings for RDSMC and all baseline methods. §˜E.2, E.3 and E.4 present supplementary information and additional results for the target distributions used in the main experiments. §˜E.5 reports additional experiments that control for hyperparameters, runtime, and other relevant factors to compare across methods, as well as an empirical comparison to phillips2024particle.